# FoundPose: Unseen Object Pose Estimation with Foundation Features

### v1

@@ Abstract
We propose FoundPose, a method for 6D pose estimation of unseen rigid objects from a single RGB image. The method assumes that 3D models of the objects are available but does not require any object-specific training. This is achieved by building upon DINOv2, a recent vision foundation model with impressive generalization capabilities. An online pose estimation stage is supported by a minimal object representation that is built during a short onboarding stage from DINOv2 patch features extracted from rendered object templates. Given a query image with an object segmentation mask, FoundPose first rapidly retrieves a handful of similarly looking templates by a DINOv2-based bag-of-words approach. Pose hypotheses are then generated from 2D-3D correspondences established by matching DINOv2 patch features between the query image and a retrieved template, and finally optimized by featuremetric refinement. The method can handle diverse objects, including challenging ones with symmetries and without any texture, and noticeably outperforms existing RGB methods for coarse pose estimation in both accuracy and speed on the standard BOP benchmark. With the featuremetric and additional MegaPose refinement, which are demonstrated complementary, the method outperforms all RGB competitors. Source code is at: evinpinar.github.io/foundpose.
@@ 1 Introduction
Image-based estimation of the 6D object pose (3D rotation and 3D translation) is a crucial research problem in the field of spatial AI. In robotics, for example, the information about object poses allows a robot to act upon the objects which enables fully automated solutions for warehouse operation or assembly. In mixed reality applications, this information unlocks physical interaction with replicas of real-world objects, such as a computer keyboard for effective text input when fully immersed.
@@
In this work, we address the problem of model-based 6D pose estimation of unseen objects. We assume that 3D models of the target objects are available and that budget for onboarding the objects is limited (e.g., not sufficient for rendering a large-scale dataset and training a neural network). This is a practical problem setup for many applications since efficient object onboarding is often a key requirement and 3D object models can be obtained from the manufacturer or readily reconstructed [59, 76, 99].
Figure 1: Standing on the shoulders of DINOv2. Patch features from the DINOv2 foundation model (top) are the key enabler of FoundPose. Without any training, a new object is onboarded just by rendering simple templates showing the object in different orientations, and by constructing a minimal feature representation that can support online pose estimation. Thanks to the generalization capability of DINOv2 features, it is possible to estimate the pose by direct feature matching between a query (left) and a retrieved template (right), despite the real-to-synthetic domain gap.
@@
The very first methods for object pose estimation can in fact handle unseen objects. These methods rely on matching hand-crafted image features [77, 50, 78] or template matching  [58, 29] and do not require any training. Later, with the rise of machine learning techniques in computer vision, most object pose estimation methods started to rely on deep neural networks. This shift brought a significant improvement in pose estimation accuracy [89] but limited generalization capability as large numbers of training images and a long training process are typically necessary for a new object instance or category. As a result, majority of these methods focus on a small set of object instances or categories. Only recently, with the accuracy scores of seen object pose estimation slowly saturating, the research field started to focus again on unseen objects, with the first attempts achieving noticeably lower accuracy scores while being computationally more demanding [82, 41, 11, 34].
@@
With their impressive generalization capabilities, foundation models [7] provide a solid ground for solving the problem at hand. Models such as DINOv2 [65, 14], CLIP[75] or ALIGN [37] have been successfully applied on various vision tasks without any fine-tuning [56, 97, 23, 60]. For example, the CNOS [60] method for segmentation of unseen objects, which leverages DINOv2 [65, 14] and Segment Anything [40], outperforms Mask R-CNN [25] without any training on the target objects.
Inspired by these success stories, we propose FoundPose, a method for model-based pose estimation of unseen objects that revisits classical computer vision techniques with the power of DINOv2, without requiring any training. Given an object segmentation mask from CNOS [60], we first rapidly retrieve a small set of similarly looking templates using bag-of-words descriptors [84], which are constructed from DINOv2 patch descriptors. For each retrieved template, we establish 2D-3D correspondences by matching DINOv2 patch descriptors of the query image and the template, and estimate a pose hypothesis by PnP-RANSAC[43, 20]. Finally, we refine the best hypothesis by featuremetric alignment. As demonstrated in our experiments, FoundPose works on diverse objects, including challenging objects with symmetries and without any texture.
In this work, we make the following contributions:
@@
    A practical method for unseen object pose estimation which demonstrates the effectiveness of DINOv2 for this task by outperforming existing methods on the BOP benchmark [34] in both accuracy and speed, without any training.
@@
    A rapid template retrieval approach which is based on visual words constructed from DINOv2 patch descriptors, is robust to partial object occlusions, and considerably simplifies the subsequent problem of establishing 2D-3D correspondences.
@@
    A minimal template-based object representation which is fast to build (requiring only several hundreds of rendered object templates), can effectively support online pose estimation, and has a low storage footprint enabling to scale to large numbers of objects (only bag-of-words descriptors and patch descriptors registered in 3D are stored).
@@
    A featuremetric pose refinement approach which is effective in reducing discrepancy in 2D-3D correspondences due to coarse patch sampling, and complementary to the MegaPose refiner (achieving more together than any of them alone).
@@ 2 Related work
This paper builds on over 60 years of research in object pose estimation, and on the recent large-scale vision foundation models.
Figure 2: FoundPose overview. During a short onboarding stage, we render RGB-D templates showing the object in different orientations, extract DINOv2 patch descriptors [65, 14] from the RGB channels and register the descriptors in 3D using the depth channel. At inference time, we crop the RGB query image around the object mask predicted by CNOS [60] and retrieve a small set of most similar templates using a bag-of-words approach (with words defined by k-means clusters of patch descriptors from all templates). For each retrieved template, a pose hypothesis is generated by PnP-RANSAC [20, 43] from 2D-3D correspondences established by matching DINOv2 patch descriptors of the image crop and the template. Finally, the pose hypothesis with the highest number of inlier correspondences is refined by featuremetric alignment and reported as the final pose estimate.
Pose estimation of rigid objects. Estimating the 6D pose of a rigid object with a known 3D model from a single image is one of the oldest computer vision problems [77]. It was originally addressed by establishing 2D-3D correspondences using locally invariant features [77, 50, 78, 13], or by template matching using a global descriptor of the object appearance [58, 29]. Several techniques to build local and global descriptors have been proposed in the literature. The first descriptors were manually designed and relied on intensity edges [49, 36, 19, 93], raw pixels [58, 78, 6, 55, 64] or image gradients [51, 29, 28]. Methods based on hand-crafted features have been progressively replaced and outperformed [32, 33, 34] by deep neural networks [98, 91, 90, 68, 102, 85, 30, 46, 35, 69, 70]. State of the art methods [89, 34, 96] typically do not rely on explicit feature representation and instead directly regress either the absolute [39, 100, 96] or relative pose [53, 45, 42, 41]. Methods based on direct absolute pose regression are trained for specific object instances or categories and do not generalize to unseen objects. Methods based on direct relative pose estimation usually rely on render-and-compare [45], which requires large amounts of synthetic images showing diverse objects to achieve generalization capabilities [41]. We build on earlier works on template matching and correspondence estimation, but use features extracted with a pre-trained DINOv2 model [65] instead of hand-crafted features. Our method does not require any training and can be directly applied to any object.
Pose estimation of unseen rigid objects. Motivated by practical applications, several methods that can be deployed to new objects without re-training have been proposed. Some of the methods address unseen object pose estimation without a 3D model [27, 47, 87, 26] by capturing multiview reference views and relative poses.Others propose learning local or global descriptors that can generalize to unseen objects. For example [98, 61, 82, 88, 92] learn global descriptors for template matching using contrastive learning [98, 61, 82, 92] or an auto-encoder [88]. A sparse set of generic 3D keypoints is predicted in[74, 73][82] uses a network to predict dense correspondences, and  [101] directly predicts the 3D object orientation. Other methods rely on depth measurements [67, 4, 63, 11]. In this work, we consider the more challenging scenario where only RGB images are available.
The state of the art [34] methods [41, 57] for 6D pose estimation of unseen objects rely on a three-stage approach as follows (1) 2D detection using CNOS [60], (2) a coarse pose is estimated by a neural network that predicts similarity between templates and the cropped input image, and (3) pose refinement by a regression network based on the render-and-compare strategy [62, 45, 42, 57, 24]. These methods achieve generalization to novel objects by training on millions of images showing thousands of different objects. Such datasets [41] can only be obtained in simulation, which exposes them to the synthetic-to-real domain gap. Our method leverages a foundation model, and without any re-training, significantly outperforms previous coarse pose estimation strategies in terms of accuracy and runtime, and achieves the state of the art when combined with the refinement network of MegaPose [41]. Concurrent to our work is ZS6D [3] which also uses a foundation model for pose estimation and to which we compare in Sec. 4.
@@
Foundation models on related tasks. A foundation model is a machine-learning model trained on broad data using self-supervised learning that can be adopted to a wide range of downstream tasks [7]. Foundation models initially appeared in Natural Language Processing (NLP) with examples such as BERT [15] or GPT-3[9]. In computer vision, foundation models include CLIP [75, 12], ALIGN [37], iBOT [104], ImageBind [22], DINOv1 [10], and DINOv2 [65]. The amount of data [80, 65] used to train these models and their performance [65] on downstream tasks are constantly growing. The models have achieved results on-par or better than supervised models for several fundamental computer vision tasks, such as image classification [10, 12, 65], semantic segmentation [56], object discovery [97, 83, 66], and depth estimation [65]. DINOv2, which is based on a Vision Transformer architecture[16] and trained in self-distillation fashion, has been successfully employed in several works. Thanks to the attention architecture which meets the global context with local image features, it has been shown to encode both the high spatial information about the object parts as well as semantic information such as object categories [2]. It has been successfully used in zero-shot setups, i.e., without any training, for establishing semantic correspondences [2, 103, 23, 48]. Building on top of these recent insights about foundation models, we present an approach that uses DINOv2 for the task of 6D pose estimation of unseen objects and improves the state of the art.
@@ 3 FoundPose
In this section we describe FoundPose, the proposed method for unseen object pose estimation. We first provide a high-level overview of the method in Sec. 3.1 and then focus on the key components and rationale of our design choices in Sec. 3.23.5.
@@ 3.1 Method overview
Problem definition. We consider the problem of estimating the 6D pose of rigid objects from a single RGB image with known intrinsics. The objective is to estimate the pose of all instances of target objects that are visible in the image. We assume that the only information provided for the target objects are their 3D mesh models, and there is only a limited budget for onboarding the objects, i.e., for preparing object representations that can be used for pose estimation. We assume that segmentation masks of the target object instances, together with per-mask object identity, are provided at inference time. In our experiments, we obtain the masks by CNOS [60], a recent method for segmentation of unseen objects that similarly is based on DINOv2 and requires only 3D models for onboarding the objects.
Onboarding and inference. During an offline onboarding stage, we render templates showing 3D object models in different orientations. From each template, we extract DINOv2 descriptors of image patches and register the descriptors in 3D, i.e., each patch descriptor is associated with the corresponding 3D location in the object model space (Sec. 3.2). At inference time, given a segmentation mask of an object instance, we crop the image region around the mask, extract DINOv2 patch descriptors of the crop, and use them with a bag-of-words retrieval technique to efficiently identify a small set of templates that show the object in orientations similar to the observation (Sec. 3.3). For each retrieved template, we establish 2D-3D correspondences by matching DINOv2 patch descriptors from the crop against patch descriptors of the templates registered in 3D and generate a pose hypothesis using the PnP-RANSAC algorithm (Sec. 3.4). Finally, we refine the best pose hypothesis by featuremetric alignment, an optimization-based algorithm inspired by photometric alignment that operates on features (Sec. 3.5). The pipeline of the method is shown in Fig. 2.

### v2

@@ Abstract
We propose FoundPose, a model-based method for 6D pose estimation of unseen objects from a single RGB image. The method can quickly onboard new objects using their 3D models without requiring any object- or task-specific training. In contrast, existing methods typically pre-train on large-scale, task-specific datasets in order to generalize to new objects and to bridge the image-to-model domain gap. We demonstrate that such generalization capabilities can be observed in a recent vision foundation model trained in a self-supervised manner. Specifically, our method estimates the object pose from image-to-model 2D-3D correspondences, which are established by matching patch descriptors from the recent DINOv2 model between the image and pre-rendered object templates. We find that reliable correspondences can be established by kNN matching of patch descriptors from an intermediate DINOv2 layer. Such descriptors carry stronger positional information than descriptors from the last layer, and we show their importance when semantic information is ambiguous due to object symmetries or a lack of texture. To avoid establishing correspondences against all object templates, we develop an efficient template retrieval approach that integrates the patch descriptors into the bag-of-words representation and can promptly propose a handful of similarly looking templates. Additionally, we apply featuremetric alignment to compensate for discrepancies in the 2D-3D correspondences caused by coarse patch sampling. The resulting method noticeably outperforms existing RGB methods for refinement-free pose estimation on the standard BOP benchmark with seven diverse datasets and can be seamlessly combined with an existing render-and-compare refinement method to achieve RGB-only state-of-the-art results. Project page: evinpinar.github.io/foundpose.
footnotetext: *Work done during Evin’s internship at Meta.
@@ 1 Introduction
Figure 1: Bridging synthetic-to-real gap. Patch descriptors from an intermediate layer of DINOv2[62] (top), a recent vision foundation model, are the key enabler of FoundPose. Thanks to the generalization capability of these descriptors, it is possible to establish reliable correspondences between a real query image (left) and a synthetic template (right) by a simple nearest-neighbor matching. The patch descriptors are colored by the top three components of a PCA space computed from descriptors of all object templates. Note that colors of the same object parts are consistent, despite the real-to-synthetic domain gap.
@@
Image-based estimation of the 6D object pose (3D rotation and 3D translation) is an important research problem in the field of spatial AI. In robotics, for example, the information about object poses allows a robot to act upon the objects, which enables fully automated solutions for warehouse operation or assembly. In mixed-reality applications, this information unlocks physical interaction with replicas of real-world objects, such as a computer keyboard, for effective text input when fully immersed.
@@
In this work, we address the problem of model-based 6D pose estimation of unseen objects. We assume that 3D models of the objects are available and that budget for onboarding the objects is limited (e.g., not sufficient for rendering a large-scale dataset and training a neural network). This is a practical problem setup for many applications since efficient object onboarding is often a key requirement and 3D object models can be obtained from the manufacturer or readily reconstructed [57, 71, 90].
@@
The very first methods for object pose estimation can, in fact, handle unseen objects as they do not require any training, typically just a set of reference images. These methods rely on classical techniques such as matching hand-crafted image features [72, 49, 69, 14] or template matching  [55, 29]. Later, with the rise of machine learning techniques in computer vision, most object pose estimation methods started to rely on deep neural networks. This shift brought a significant improvement in pose estimation accuracy [82] but limited generalization capability as large numbers of training images and a lengthy training process are usually necessary for every new object instance or category. As a result, the majority of these methods focus on a small set of objects. Only recently, with the accuracy scores of seen object pose estimation slowly saturating, the research field started to focus again on unseen objects [56], with the first attempts achieving noticeably lower accuracy scores while being computationally more demanding [75, 41, 12, 34].
@@
With their impressive generalization capabilities, foundation models [7] provide a solid ground for solving the problem at hand. Models such as DINOv2 [62, 15], CLIP[70] or ALIGN [37] have been successfully applied on various vision tasks without any task-specific training [53, 88, 24, 58]. For example, CNOS [58] leverages frozen DINOv2 [62, 15] and Segment Anything [40] and outperforms Mask R-CNN [25] on the object segmentation task. Goodwin et al. [24] show that DINO patch descriptors[11] can be used to establish semantic correspondences between instances of the same object category.
Inspired by these success stories, we propose FoundPose, a method for model-based pose estimation of unseen objects, which brings the power of modern foundation features into classical computer vision techniques via careful design choices. Despite being surprisingly simple, easy-to-interpret, and requiring no object- or task-specific training, the method achieves state-of-the-art results on the standard BOP benchmark [34].
First, given an RGB image and an object mask from CNOS [58], we perspectively crop the object region and rapidly retrieve a small set of similarly looking, pre-rendered object templates. To this end, we develop an efficient retrieval approach by integrating DINOv2 into the bag-of-words representation from 2003 [76]. This approach is 15X faster while only slightly less accurate than the heavy render-and-compare coarse stage of MegaPose [41], and requires 100X less templates (several hundreds vs 90K+) than previous approaches [75, 60] (the overall memory footprint is 25X lower).
Second, we establish 2D-2D correspondences between each retrieved synthetic template (with fixed light and black background) and the real image crop by simple one-way kNN search of DINOv2 patch descriptors (Fig. 1). In contrast, existing methods typically train on large-scale, heavily randomized, and task-specific datasets to bridge the synthetic-to-real domain gap [33, 41, 83]. We demonstrate that patch descriptors from an intermediate DINOv2 layer, which were shown to carry stronger positional information [2], are crucial for achieving geometrically consistent correspondences when semantic information is ambiguous due to object symmetries or a lack of texture. We show that the intermediate DINOv2 descriptors are in fact the key enabler of FoundPose, yielding significantly higher accuracy also compared to descriptors extracted with SAM[40], CLIP [70], LoFTR [78], S2DNet [23], and dense SIFT [49]. Next, for each retrieved template, we generate a pose hypothesis from image-to-model 2D-3D correspondences, which are established by lifting the matched 2D patch locations in the template to 3D using rendered depth. Finally, we further optimize the top-quality hypothesis by featuremetric refinement, which applies the idea of the classical photometric refinement [4] to DINOv2 patch descriptors. The refinement effectively compensates for the discrepancy in the 2D-3D correspondences caused by coarse sampling of DINOv2 patches.
In summary, we make the following contributions:
@@
    A training-free method for model-based object pose estimation which relies on a surprising simple and easy-to-interpret DINOv2-based pipeline and achieves state-of-the-art results on the standard BOP benchmark [34].
@@
    An efficient template retrieval approach which requires 100X fewer templates than previous approaches and is robust to partial object occlusions.
@@
    A lightweight object representation which is fast to build and has a 25X lower memory footprint than competitors, enabling scaling to large numbers of objects.
@@
    A featuremetric refinement approach which compensates for coarse patch sampling.
  5. 5.
    Demonstrated importance of intermediate DINOv2 descriptors for handling symmetric and texture-less objects, also outperforming descriptors from other foundation models.
@@ 2 Related work
This paper builds on over 60 years of research in object pose estimation and on the recent large-scale vision foundation models.
Classical methods. Estimating the 6D pose of rigid objects from a single image is one of the first computer vision problems [72]. Early methods relied on local feature matching [49, 69, 14] or template matching [55, 29], and could rapidly onboard new objects if provided with a set of reference images annotated with model-to-camera transformations. With the introduction of Microsoft Kinect, the attention of the research field was steered towards object pose estimation from RGB-D or D-only images, yielding methods based on 3D local features[85], notably successful point-pair features[21, 32], and methods based on RGB-D template matching[28, 36]. The RGB-D methods produce more accurate poses and are therefore popular in industry, but their application in open-world scenarios is limited. Besides boosting the pose estimation accuracy, the additional depth channel from Kinect-like sensors enabled easy 3D object reconstruction [57], and, in turn, methods relying on 3D mesh models started to emerge. The model-based object pose estimation setup is still popular [82, 34], among both RGB and RGB-D methods, and is relevant for factory and warehouse scenarios where CAD object models are often available. On the other hand, the model-free setup, recently revisited in [79, 26], is relevant for mixed reality applications where the set of target objects is typically small and capturing reference images is easy.
Deep learning methods. As in other fields of computer vision, methods based on hand-crafted features and techniques have been progressively replaced by methods based on deep neural networks[91, 51, 45, 42, 87], which can operate on RGB or RGB-D inputs. These methods represent the current state of the art in terms of accuracy [82, 34]. However, their scalability is hindered by the requirement of a large-scale training dataset for learning new objects. To address this issue, deep-learning methods that can onboard new objects without any object-specific training have been proposed recently. As examples of model-based methods, Nguyen et al. [60], Shugurov et al. (OSOP)[75], and Thalhammer et al. [84] learn descriptors for template matching by contrastive learning, Sundermeyer et al. [80] generate such descriptors by an augmented auto-encoder, Pitteri et al. [68, 67] predict generic 3D keypoints or local surface embeddings, and Xiao et al. [92] directly predict the 3D object orientation. Model-free methods [27, 48, 79, 26], and methods relying on depth measurements [64, 5, 61, 12] have been also proposed.
The top-performing methods on the unseen object pose estimation task of the BOP Challenge 2023 [34] include GenFlow[54], MegaPose[41], GigaPose[59], FoundationPose[89], SAM-6D[46], and PoMZ[10]. To achieve generalization to novel objects, all except PoMZ require generating millions of task-specific training images showing thousands of different objects. Generating such datasets requires significant effort and opens up new types of challenges, including positioning objects in the scene[35, 16], collecting a sufficiently large set of object models, or texturing the models[89]. In contrast, FoundPose does not require any training, uses frozen DINOv2 features, and outperforms RGB-only GenFlow, MegaPose, and GigaPose. FoundationPose, SAM-6D, PoMZ, as well as ZeroPose[12], are RGB-D methods and do not have RGB-only versions. To the best of our knowledge, the only training-free methods for model-based object pose estimation are PoMZ[10], which requires RGB-D inputs, and ZS6D[3], which achieves significantly lower accuracy than FoundPose. ZS6D uses features from the last ViT [18] layer for establishing correspondences, which we show inferior to our solution (Sec. 4). Note that [54, 12, 3, 10, 59, 89] are all unpublished at the time of submission.
@@
Foundation models. A foundation model is a machine-learning model trained on broad data by self-supervised learning that can be adapted to a wide range of downstream tasks [7]. Foundation models initially appeared in natural language processing with examples such as BERT [17] or GPT-3[9]. In computer vision, foundation models already achieve on-par or better results than supervised models [7, 11, 62, 13, 88, 53, 63]. A prominent example is DINOv2 [62], which is based on the Vision Transformer architecture[18], trained in a self-distillation fashion, and has been shown to encode fine spatial information about the object parts as well as semantic information about object categories [2]. It has been successfully used in zero-shot setups, i.e., without any training, for establishing semantic correspondences [2, 93, 24, 47]. FoundPose builds on these insights. Compared to Goodwin et al. [24] which use DINO patch descriptors to establish semantic correspondences within an object category, FoundPose shows that DINOv2 descriptors can be used to establish synthetic-to-real correspondences. Furthermore, [24] requires RGB-D inputs at test time, cannot handle symmetric objects (such objects are omitted in their evaluation), and is compared only with custom baselines. FoundPose assumes RGB-only test inputs, can handle symmetric objects by design, and achieves state-of-the-art results on the standard BOP benchmark. From the already reviewed methods for pose estimation of specific objects, PoMZ[10] and ZS6D[3] also rely on frozen DINOv1/v2. However, the first needs RGB-D inputs, and the latter achieves noticeably lower accuracy.
@@ 3 FoundPose
In this section, we describe FoundPose, the proposed method for unseen object pose estimation. We first provide a high-level overview of the method in Sec. 3.1 and then focus on the key components and rationale of our design choices in Sec. 3.23.5.
@@ 3.1 Method overview
Problem definition. We consider the problem of estimating the 6D pose of rigid objects from a single RGB image with known intrinsics. The objective is to estimate the pose of all instances of target objects that are visible in the image. We assume that the only information provided for the target objects are their 3D mesh models and that there is only a limited budget for onboarding the objects, i.e., for preparing object representations that can be used for online pose estimation. We constrain the onboarding process to 5 minutes and 1 GPU, as required by the BOP Challenge 2023[34]. We additionally assume that segmentation masks of the target object instances, together with per-mask object identity, are provided at inference time. In our experiments, we obtain the masks by CNOS [58], a recent method for segmentation of unseen objects that also requires only 3D models for onboarding the objects.
Onboarding and inference. During an offline onboarding stage, we render templates showing 3D object models in different orientations. From each template, we extract DINOv2 descriptors of image patches and register the descriptors in 3D, i.e., each patch descriptor is associated with the corresponding 3D location in the object model space (Sec. 3.2). At inference time, given a segmentation mask of an object instance, we crop the image region around the mask, extract DINOv2 patch descriptors of the crop, and apply a bag-of-words retrieval technique to efficiently identify a small set of templates that show the object in orientations similar to the observation (Sec. 3.3). For each retrieved template, we establish 2D-3D correspondences by matching patch descriptors from the crop against patch descriptors from the template, and generate a pose hypothesis by the PnP-RANSAC algorithm (Sec. 3.4). Finally, we refine the best pose hypothesis by featuremetric alignment, an optimization-based algorithm inspired by photometric alignment that operates on features (Sec. 3.5). The pipeline of the method is shown in Fig. 2.
Figure 2: FoundPose overview. During a short onboarding stage, we render RGB-D templates showing the object in different orientations, extract DINOv2 patch descriptors [62, 15] from the RGB channels and register the descriptors in 3D using the depth channel. At inference time, we crop the RGB query image around the object mask predicted by CNOS [58] and retrieve a small set of most similar templates using a bag-of-words approach (with words defined by k-means clusters of patch descriptors from all templates). For each retrieved template, a pose hypothesis is generated by PnP-RANSAC [22, 43] from 2D-3D correspondences established by matching patch descriptors of the image crop and the template. Finally, the pose hypothesis with the highest number of inlier correspondences is refined by featuremetric alignment.